log(2) (x^2 - 1) = 3
Answer by nerdybill(7384)

log(2) (x^2 - 1) = 3
(x^2 - 1) = 2^3
(x^2 - 1) = 8
x^2 = 9
x = (+-)3
.
that's plus or minus 3